# AI for Tabular Data

### v1

@@ B.1 Outlier Detection
[…]
B.2 Classification and Regression
The AUC distribution for binary and multiclass classification and the MAPE distribution for regression for the three investigated embedding models and the two different algorithms (KNN, MLP) are shown in Figures 2, 3 and 4. Since the number of suited TabArena datasets was limited at the time of this investigation, boxplots showing the distribution of the results may not be expressive, especially for multiclass classification and regression, where only 6 datasets where available. As a continuously updated benchmark for machine learning on tabular data, TabArena is expected to expand its repository with increasingly appropriate datasets.
Figure 2: Distribution of the performance for binary classification (AUC) for the three embedding models shown in Table 2.
Figure 3: Distribution of the performance for multiclass classification (AUC) for the three embedding models shown in Table 3.
Figure 4: Distribution of the performance for regression (MAPE) for the three embedding models shown in Table 4.

### v2

@@ 4 Results
Table 1 refers to the average time needed to compute the embedding for one table row. Median, mean and standard deviation are calculated across all datasets. Note that the computation time is not directly comparable, since TabICL, TabPFN and TabSTAR use the GPU, whereas TableVectorizer and the sphere model run on the CPU. Nevertheless, TabICL, TableVectorizer and the sphere models are faster than TabPFN and TabSTAR. TabPFNColumns computes the embedding for each column and aggregates them by averaging, hence it is clear that TabPFNColumns requires significantly more computation time.
   Method Median (ms) Mean (ms) Std (ms)
   TableVectorizer  0.17  0.24  0.28
   Sphere (Dim 192)  0.08  0.16  0.21
   Sphere (Dim 256)  0.09  0.18  0.25
   Sphere (Dim 512)  0.17  0.32  0.45
   TabICL  0.16  0.35  0.50
   TabPFNColumns 28.06 153.61 528.91
   TabPFNConstant  1.49  2.96  7.44
   TabPFNRandom  1.97  4.15  7.73
   TabSTAR  1.06  1.93  3.34
Table 1: Average embedding time for one table row (one sample), compiled over all datasets.
Figure 12: Critical difference diagram for supervised tasks.
Figure 13: Critical difference diagram combined for all tasks.
5 Discussion
Overall, Fig. 13 presents a critical difference diagram combining every evaluation across all datasets. The diagram shows that TableVectorizer and the sphere models perform significantly better than TabICL and the TabPFN versions. Furthermore, TableVectorizer and TabSTAR outperform TabPFN. In summary, on the investigated benchmarks, classical feature engineering methods perform better than task-agnostic embeddings extracted from tabular foundation models. However, general ranking conclusions for the investigated embedding models that go beyond the evaluated datasets cannot be drawn from these results, as the number of datasets is limited. Furthermore, we extracted task-agnostic embeddings as described in Section 2. There might be ways to improve the extraction.
Regarding the embedding dimensionality, TableVectorizer constructs embeddings with the lowest dimension ($\leq$ 166), while TabPFN requires 192, TabSTAR 384 and TabICL even 512 as embedding dimension. A significant limitation of TableVectorizer is its variable embedding dimensionality, which is determined by the input dataset. This dataset-dependent dimensionality poses a challenge for applications requiring the joint embedding of diverse datasets into a unified representation space. In contrast, models like TabPFN and TabICL employ a fixed, but significantly larger embedding size across all datasets. The advantage of the simple sphere model is its arbitrary embedding dimension, while keeping a competitive or even superior performance. The choice of the dimension, e.g., 192, 256 or 512, does not seem to significantly influence the performance.
Compared to original end-to-end tabular foundation models TabPFNOriginal and TabICLOriginal, the best task-agnostic embedding models (followed by KNN or MLP) do not exhibit significant performance loss. Moreover, the task-agnostic embeddings are reusable and could be applied in unsupervised settings such as similarity search, where the original end-to-end tabular foundation models cannot be applied a priori. On the other hand, these end-to-end models provide a direct prediction, while prediction based on task-agnostic embeddings may require intense hyperparameter training.
6 Conclusion and Future Work
This study demonstrates that simple feature engineering methods like TableVectorizer or a simple sphere model achieve competitive performance for task-agnostic embeddings, while requiring substantially fewer computational resources. Prediction results based on task-agnostic embeddings are not significantly worse even in settings where sophisticated tabular foundation models can fully leverage their in-context learning capabilities.
However, several open challenges remain for future work. First of all, our findings have to be confirmed on a wider range of datasets and more embedding models, e.g., TabPFNv2.5 [5], Mitra [13] and LimiX [12]. Second, although the embeddings are constructed in a task-independent way, their task-agnostic quality should be further assessed considering additional downstream tasks and metrics.
Furthermore, current methods primarily capture statistical patterns. Recent works such as ConTextTab [11] and TabSTAR [1] demonstrate promising directions by incorporating textual (meta-)data. A comprehensive analysis for their task-agnostic embeddings incorporating text would be a further step toward multi-modal alignment, enabling tabular representations to be aligned with text or images for cross-modal retrieval and reasoning. The introduced sphere model has an advantage for aligning with other modalities, due to its arbitrary embedding dimension.
Large amounts of data occur subsequently and not as a closed dataset. The ingestion of new rows raises fundamental questions about context selection for in-context learners: Should entire datasets be re-embedded, or can a representative subset suffice? Distribution shift poses additional complications. The behavior of (in-context) embedding models needs to be investigated in such scenarios.
Extension to relational databases represents a natural progression, where tabular embedding models could serve as building blocks for relational models that learn representations across joined tables and foreign-key relationships. Shared representation spaces across diverse datasets are also worth investigation: Do models sufficiently differentiate between datasets when embedded jointly, or does cross-dataset contamination occur?
Addressing all these challenges will be essential for understanding (task-agnostic) tabular embeddings as a flexible, reusable tool in the broader machine learning ecosystem.